# floating point operation IVP initial value problem

### v1

@@ V-B Graph representation
Figure 1: Steps to turn power grid into $\mathcal{G}_{sim}$. First, map each component, e.g. generator and lines, to an edge $e$ whose weight $\hat{c}_{e}$ represents the FLOPs needed to solve for the current $i_{e}$. Second, build a graph that resembles the grid topology.
Let $\mathcal{G}_{grid}=(\mathcal{V}_{grid},\mathcal{E}_{grid})$ be a graph representing the grid topology, that is, the nodes $n\in\mathcal{V}_{grid}$ represent buses and the edges $e\in\mathcal{E}_{grid}$ represent the electrical transmission lines. To put all the components, i.  e. generators and transmission lines, on an equal footing, we construct a simulation graph $\mathcal{G}_{sim}=(\mathcal{V}_{sim},\mathcal{E}_{sim})\supset\mathcal{G}_{grid}$ based on the grid $\mathcal{G}_{grid}$ in the following way: Except the transmission lines that are already represented as edges, we represent each component attached to a bus $n\in\mathcal{V}_{grid}$ by an additionally added edge [^1] We also add an additional node for each component. that is incident to the bus $n$. Hence all electrical components are represented by edges in the graph $\mathcal{G}_{sim}$. To all edges, we then assign a weight $\hat{c}_{e}=c_{e}/c_{max}\in[0,1]$ that represents the normalized computational cost (normalized FLOPs) of simulating the respective component, cf. Section V-A. The steps to construct $\mathcal{G}_{sim}$ are illustrated in Fig. 1.
The problem of finding an optimal partitioning of the network into two sub-networks for parallel simulation is thus equivalent to finding an optimal partitioning of the graph $\mathcal{G}_{grid}$ taking into account all components whose computational costs are represented by the weights $c_{e}$ of the edges in $\mathcal{G}_{sim}$. We only need to partition the graph $\mathcal{G}_{grid}\subset\mathcal{G}_{sim}$ as the nodes in $\mathcal{V}_{sim}\setminus\mathcal{V}_{grid}$ don’t represent any additional partition choices regarding the network and thus the components: All edges representing components between nodes in different parts $\mathcal{V}_{i}$ are in the cut set. For all other components, there is at least one node $n\in\mathcal{V}_{grid}$ incident whose assignment to $\mathcal{V}_{1}$ or $\mathcal{V}_{2}$ provides the assignment of the component to part 1 or part 2.
@@ V-C QUBO Formulation
$Q_{part.}=Q_{components}+Q_{cut}+Q_{network}.$ (20)
Due to this construction, we avoid the additional problem of choosing the weights $\lambda_{i}$ a priori.
V-C1 Reduce ’idle’ time in component simulation
For the component simulation step, we want to minimize the idle time that arises if one part can be simulated faster than the other, see Eq. (17). By adopting from $Q_{weights}$ in Eq. (10), a QUBO objective function $Q_{components}$ that computes the squared idle time for a given partitioning is given by
$\begin{split}Q_{components}&=\sum_{n,m\in\mathcal{V}_{grid}}z_{n}\left(4\alpha(n)\alpha(m)\right)z_{m}\\ &\quad+\sum_{n\in\mathcal{V}_{grid}}z_{n}\left(-4\alpha(n)\beta\right)z_{n}+\beta^{2},\end{split}$ (21)
where $\alpha(n):=\sum_{m\in\mathcal{V}_{sim}}A_{n,m}(\mathcal{G}_{sim})\hat{c}_{n,m}$ and $\beta=\sum_{m\in\mathcal{V}_{grid}}\alpha(m)$.
@@ V-C2 Minimize the computational costs of cut simulation
for the edges in $\mathcal{G}_{grid}$. We normalize the costs attributed to the network and the injections by $c_{max}$ to obtain normalized flops.
V-C3 Reduce ’idle’ time in network simulation
@@
$Q_{network}=\frac{(2N-1)^{2}}{c_{max}^{2}}Q_{size}$ (23)
@@ V-D Partitioning for more than two computational nodes
Suppose we seek to partition the grid into $P$ parts. In that case, a direct approach is to define integer variables $z_{n}\in\{0,1,\ldots P-1\}$ to encode the partitioning and then extend the QUBO objectives to the more general case. However, encoding integer variables into binary variables significantly increases the complexity, in particular, the number of qubits. Hence, even small problem instances become impossible to embed and solve on near-term hardware.
@@ V-E Ensuring Connectedness
Ensuring connectedness is a similar problem to the traveling salesperson problem (TSP), where a satisfactory solution describes a single tour passing through all the cities that need to be visited. A known issue in the Dantzig-Fulkerson-Johnson formulation [49] of the TSP is that solutions often contain sub-tours, i.e., tours through subsets of cities. An approach for dealing with these disconnected sub-tours is the ”finger-in-the-dike” method [50], where the problem is initially solved without banning sub-tours, and ad-hoc constraints are introduced iteratively to the original formulation to ban them as they appear in the sampled solutions. Analogously, here, one could ban disconnected sub-networks as they are sampled to ensure the connectedness of the sub-networks.
@@ VI-A Optimal Solutions
Figure 2: Optimal partitionings for Case 30 (left) and a modified Case 30 (right), where one generator has been moved to a different bus. The optimal solutions have been obtained using Gurobi.
We use Gurobi to find the optimal solutions of $Q_{part.}$ for the test cases with 14, 30 and 57 buses. The MIP gap is set to 0.01. Obtaining the optimal solutions using a deterministic solver is necessary to evaluate the quality of the lowest energy states sampled by the D-Wave annealer.
The optimal solutions for a small test case demonstrate that minimizing $Q_{part.}$ gives a reasonable partition into two sub-networks, for example, cf. Fig. 2 for the 30-bus system. The sub-network sizes and the components are balanced and the cut sets are small. Furthermore, changes or modifications in the network, e.g., moving a generator, can lead to a different partitioning; see the right panel in Fig. 2. Changing the position of a generator is clear not a real life scenario but network reconfigurations are, a different optimal partitioning exists for each configuration. However, the partitioning itself introduces an overhead: $Q_{part.}$ has to be minimized, and LU decomposition has to be performed. Therefore, if a change is only persistent for a short time, re-partitioning may increase the overall computational cost of the simulation. In our future work, we plan to investigate whether we can include the overhead of repartitioning in the framework.
@@ VI-B Scaling of Computational Resources
By construction, the number of logical variables $N_{q,log}$ is given by the number of buses $N$ in the grid. $Q_{components}$ (21) and $Q_{network}$ (23) are fully connected QUBO objectives, i.e., each logical qubit has a nonvanishing coupling to every other logical qubit. Thus the full QUBO has $N_{c,log}=N^{2}/2$ logical interactions between logical qubits.
Since complete graphs, and thus fully connected QUBOs, can only be directly embedded for $N\leq 4$ on the D-Wave Advantage System 5.4 (JUPSI) with Pegasus topology [52, 53], the actual resources exceed the logical ones as we have to find minor embeddings for all test cases. We heuristically sample 10 minor embeddings for each test case to generate some minimal statistics. For the 200 and 300 bus test cases no embedding could be found in a reasonable time. For the smaller test cases, we find that the number of physical qubits $N_{q,phy.}$ approximately follows a power law $N_{q,phy.}\approx 0.11N^{2.04}$ in the number of logical qubits $N$. The power law scales similarly to the number of logical interactions $N_{c,log}$. Likewise, the number of physical couplers $N_{c,phy.}$ needed to embed the problem follows a similar power law $N_{c,phy.}\approx 0.56N_{c,log}^{2.09}$ in the number of logical interactions, see Fig. 3.
@@ VI-C Dependence of Solution Quality on Annealing Hyperparameters
Figure 4: Dependence of solution quality on quantum annealing hyperparameters. a: The left panel shows the dependence of the $TTS$ for $p_{S}=0.99$ and $L_{min}$ on the $RCS$ for only the best embeddings for cases 14, 30 and 57. For case 14, the right panel shows the variance in $TTS$ between different embeddings (encoded by colors). b: Dependence of the $TTS$ together with $L$ on the annealing time $T_{A}$ for the optimal embedding and chain strength combination found in a.
The quality of the sampled solution on the QPU depends strongly on the annealing hyperparameters. We evaluate the quality of a sample set by considering the time to the best solution $TTS$(5) for $p_{S}=0.99$ together with the lowest relative error $L_{min}$ (4) of the sample set. No optimal solution is obtained if $L_{min}>0$. To generate robust statistics, the sample set for a hyperparameter configuration for case 14 consists of $N_{s}=10^{4}$ samples, for case 30 and case 57 of $N_{s}=10^{5}$ samples.
Similar to the observations made by Gonzalez Calaza et al. [54], we find that the solution quality is sensitive to the used embedding and the chain strength $CS$, that is, the coupling strength of the physical qubits in a chain, see Fig. 4. For this analysis, we set the annealing time to $T_{A}=50\mu s$. Furthermore, for better comparison between the test cases, we consider the relative chain strength [54] given by
@@
For case 14, the fastest $TTS$ over all ten embeddings is given for $RCS$ between $0.1$ and $0.3$; see the right panel in Fig. 4a. Each embedding is assigned a color for better comparison. The performance of the embeddings varies significantly, especially for small $RCS<0.1$. The $TTS$ roughly follows a convex function for all embeddings. If the chain strength is too small, the chains break regularly and thus do not uniquely represent a logical qubit. On the other hand, if the chain strength is too high, the dynamics during annealing are dominated by the internal dynamics of the chains.
For the best-performing embedding for each $RCS$, the optimal solution could be found over almost the whole range of chain strengths, except for one outlier, see the left panel of Fig. 4a. The outlier with $L_{min}>0$ has a lower $TTS$, since the best sampled near-optimal solutions occur more frequently in the sample set than the optimal solutions for the other embeddings. For case 30, the $TTS$ and $L_{min}$ show similar behavior for the best embedding and chain strength combinations. The $TTS$ follows a convex function in the $RCS$ with a minimum between $0.02$ and $0.1$. However, the $TTS$ is at least 10 times slower than for case 14. For case 57, the current hardware limits are reached, at least as a global optimal solver. No optimal solutions could be found for any embedding and chain strength. Additionally (not shown in the figure), we sampled the QPU $10^{6}$ times for the best-performing combination of $RCS$ and embedding for case 57, which did not significantly improve the solution quality.
Another hyperparameter that, by definition, affects the $TTS$ is the annealing time $T_{A}$. For the best embedding and chain strength combinations from the previous analysis, the $TTS$ monotonically increases in $T_{A}$ for all three cases; see Fig. 4b. Here, we set $N_{s}=10^{5}$ for all configurations. The smaller problems, which do not seem to need hyperparameter optimization, could profit from using the fast annealing feature, which just became available, to achieve an even smaller TTS.
@@
We conclude that the performance of QA depends strongly on hyperparameters, particularly on the embedding and the chain strength. Furthermore, for quantum annealing, the actual time to solution suffers from overheads, i. e. the time needed to embed the problem and network latency. For a fair comparison between the performances of quantum annealing on current prototypical hardware with classical deterministic or heuristic solvers, the overhead must be included, and the question arises whether to compare optimized solvers. A fair comparison would thus reveal current hardware limitations of the D-Wave machine, in particular, the need for minor embeddings [56], and is thus not meaningful for the future where some problems might have been overcome, e.g., by adapting the driver to expand the minimal gap [57]. In this work, we thus omit a benchmarking study and refer to the recent literature: For other highly connected Ising problems, quantum annealing on current D-Wave machines does not provide a speed up or better solution quality [56, 58, 59].

### v2

@@ V-B Graph representation
Figure 1: Steps to turn power grid into $\mathcal{G}_{comp}$. First, map each component, e.g. generator and lines, to an edge $e$ whose weight $\hat{c}_{e}$ represents the FLOPs needed to solve for the current $i_{e}$. Second, build a graph that resembles the grid topology.
Let $\mathcal{G}_{grid}=(\mathcal{V}_{grid},\mathcal{E}_{grid})$ be a graph representing the grid topology, that is, the nodes $n\in\mathcal{V}_{grid}$ represent buses and the edges $e\in\mathcal{E}_{grid}$ represent the electrical transmission lines. To put all the components, i.  e. generators and transmission lines, on an equal footing, we construct a component graph $\mathcal{G}_{comp}=(\mathcal{V}_{comp},\mathcal{E}_{comp})\supset\mathcal{G}_{grid}$ based on the grid $\mathcal{G}_{grid}$ in the following way: Except the transmission lines that are already represented as edges, we represent each component attached to a bus $n\in\mathcal{V}_{grid}$ by an additionally added edge [^1] We also add an additional node for each component. that is incident to the bus $n$. Hence all electrical components are represented by edges in the graph $\mathcal{G}_{comp}$. To all edges, we then assign a weight $\hat{c}_{e}=c_{e}/c_{max}\in[0,1]$ that represents the normalized computational cost (normalized FLOPs) of simulating the respective component, cf. Section V-A. The steps to construct $\mathcal{G}_{comp}$ are illustrated in Fig. 1.
The problem of finding an optimal partitioning of the network into two sub-networks for parallel simulation is thus equivalent to finding an optimal partitioning of the graph $\mathcal{G}_{grid}$ taking into account all components whose computational costs are represented by the weights $\hat{c}_{e}$ of the edges in $\mathcal{G}_{comp}$. We only need to partition the graph $\mathcal{G}_{grid}\subset\mathcal{G}_{comp}$ as the nodes in $\mathcal{V}_{comp}\setminus\mathcal{V}_{grid}$ do not represent any additional partition choices regarding the network and thus the components: All edges representing components between nodes in different parts $\mathcal{V}_{i}$ are in the cut set. For all other components, there is at least one node $n\in\mathcal{V}_{grid}$ incident whose assignment to $\mathcal{V}_{1}$ or $\mathcal{V}_{2}$ provides the assignment of the component to part 1 or part 2.
@@ V-C QUBO Formulation
$Q_{part.}=Q_{comp.}+Q_{cut}+Q_{net.}.$ (18)
Due to this construction, we avoid the additional problem of choosing the weights $\lambda_{i}$ a priori. However, rescaling the objectives, like any reformulation, can affect the solution time if heuristics are involved without changing the solution qualitatively, see discussion in section VI-E.
V-C1 Reduce “idle” time in component simulation
For the component simulation step, we want to minimize the idle time that arises if one part can be simulated faster than the other, see Eq. (15). By adopting from $Q_{weights}(p=1)$ in Eq. (10), a QUBO objective function $Q_{comp.}$ that computes the squared idle time for a given partitioning is given by
$Q_{comp.}$ $=\left(2\sum_{n\in\mathcal{V}_{grid}}z_{n}\tilde{\alpha}(n)-\sum_{n\in\mathcal{V}_{grid}}\tilde{\alpha}(n)\right)^{2}$ (19)
with
$\tilde{\alpha}(n)=\frac{1}{2}\sum_{m\in\mathcal{V}_{grid}}\hat{c}_{n,m}+\sum_{m\in\mathcal{V}_{comp.}\setminus\mathcal{V}_{grid.}}\hat{c}_{n,m}.$
The definition of $\tilde{\alpha}(n)$ takes into account edges in $\mathcal{E}_{comp.}\setminus\mathcal{E}_{grid.}$ are incident to only one node in $\mathcal{V}_{grid}$ by construction, and thus, are not counted twice. On the other hand, all edges in $\mathcal{E}_{grid.}$ are incident to two nodes in $\mathcal{V}_{grid}$, and thus, a factor of $\frac{1}{2}$ is introduced in the first term of $\tilde{\alpha}(n)$ to avoid double-counting.
@@ V-C2 Minimize the computational costs of cut simulation
for the edges in $\mathcal{G}_{grid}$. Again, $N$ is the number of buses in the grid, thus $N=\lvert\mathcal{V}_{grid}\rvert$. We normalize the costs attributed to the network and the injections by $c_{max}$ to obtain normalized flops. Then, the objective function that estimates the overhead due to the sequential simulation of the cut is given by
$\begin{split}Q_{cut}=\frac{1}{2}\sum_{n,m\in\mathcal{V}_{grid}}A_{n,m}(\mathcal{G}_{grid})\left(\hat{c}_{n,m}+4\frac{N-1}{c_{max}}\right)\\ (-2z_{n}z_{m}+z_{n}+z_{m}).\end{split}$ (21)
V-C3 Reduce “idle” time in network simulation
@@
$Q_{net.}=\frac{(2N-1)^{2}}{c_{max}^{2}}\left(2\sum_{n\in\mathcal{V}_{grid}}z_{n}-N\right)^{2}$ (22)
@@ V-D Partitioning for more than two computational nodes
Suppose we seek to partition the grid into $P$ parts. In that case, a direct approach is to define integer variables $z_{n}\in\{0,1,\ldots P-1\}$ or one-hot encoded variables $z_{n,p}$, where $z_{n,p}=1$ if and only if bus $n$ is in part $p$, to encode the partitioning and then extend the QUBO objectives to the more general case. One-hot encoding has been used in the aforementioned [23] to minimize the cuts and balance the sizes of the subgrids. However, both encodings significantly increase the complexity, in particular, the number of qubits. Hence, even small problem instances become impossible to embed and solve on near-term hardware.
@@ V-E Ensuring Connectedness
Ensuring connectedness is a similar problem to the traveling salesperson problem (TSP), where a satisfactory solution describes a single tour passing through all the cities that need to be visited. A known issue in the Dantzig-Fulkerson-Johnson formulation [55] of the TSP is that solutions often contain sub-tours, i.e., tours through subsets of cities. An approach for dealing with these disconnected sub-tours is the “finger-in-the-dike” method [56], where the problem is initially solved without banning sub-tours, and ad-hoc constraints are introduced iteratively to the original formulation to ban them as they appear in the sampled solutions. Analogously, here, one could ban disconnected sub-networks as they are sampled to ensure the connectedness of the sub-networks.
@@ VI-A Optimal Solutions
Figure 2: Optimal partitionings for Case 14, case 57, case 30 and modified case 30, where one generator has been moved to a different bus. The optimal solutions have been obtained using Gurobi.
We use Gurobi on a Lenovo ThinkPad T14 with an Intel(R) Core(TM) i5-10210U CPU (4 cores, 8 Threads) and 32 GB installed RAM to find the reference solutions of $Q_{part.}$ for the test cases with 14, 30, and 57 buses. Gurobi uses a branch and bound method to solve mixed integer problems (MIP). It is a deterministic solver that uses heuristics, e.g. for node selection and cut selection. The optimal solutions (MIPGap = 0) are found in $T_{Gur.}=0.43s$ for case 14, in $T_{Gur.}=4.66s$ for case 30 and in $T_{Gur.}=1.04s$ for case 57, see also Table I and the discussion in section VI-C for a base comparison to the performance of QA.
The optimal solutions for the small test cases 14, 30 and 57 demonstrate that minimizing $Q_{part.}$ provides a reasonable partition into two sub-networks, cf. Fig. 2. The sub-network sizes and the components are balanced and the cut sets are small. Furthermore, relocating a generator, though not a realistic scenario, demonstrates how modifications or reconfigurations of the network can result in different optimal partitionings for each configuration, see the lower right panel in Fig. 2. However, it should be noted that the partitioning itself introduces an overhead, necessitating the minimization of $Q_{part.}$ and the conduction of LU decomposition. Consequently, if a change is only persistent for a brief duration, repartitioning may result in an increase of the overall computational cost of the simulation. In our future work, we plan to investigate the possibility of including the overhead of repartitioning within the presented framework.
@@ VI-B Scaling of Computational Resources
By construction, the number of logical variables $N_{q,log}$ is given by the number of buses $N$ in the grid. $Q_{comp.}$ (19) and $Q_{net.}$ (22) are fully connected QUBO objectives, i.e., each logical qubit has a nonvanishing coupling to every other logical qubit. Thus the full QUBO has $N_{c,log}=N(N-1)/2$ logical interactions between logical qubits.
Since complete graphs, and thus fully connected QUBOs, can only be directly embedded for $N\leq 4$ on the D-Wave Advantage System 5.4 (JUPSI) with Pegasus topology [58, 59], the actual resources exceed the logical ones as we have to find minor embeddings for all test cases. We heuristically sample 10 minor embeddings for each test case to generate some minimal statistics. For the 200 and 300 bus test cases no embedding could be found in a reasonable time. For the smaller test cases, we find that the number of physical qubits $N_{q,phy.}$ approximately follows a power law $N_{q,phy.}\approx 0.13N^{2.01}$ in the number of logical qubits $N$. The power law scales similarly to the number of logical interactions $N_{c,log}$. Likewise, the number of physical couplers $N_{c,phy.}$ needed to embed the problem follows a similar power law $N_{c,phy.}\approx 0.63N_{c,log}^{2.07}$ in the number of logical interactions, see Fig. 3.
@@ VI-C Dependence of Solution Quality on Annealing Hyperparameters
Figure 4: Dependence of solution quality on quantum annealing hyperparameters. a: The three panels show the dependence of the $TTS$ for $p_{S}=0.99$ and $L_{min}$ on the $RCS$ and the different embeddings for cases 14, 30 and 57. If multiple embeddings perform equal in terms of the $TTS$, only the best according to $L_{min}$ is shown here. The full data is shown Fig. 7 in the Appendix. The best hyperparameter configurations with $L_{min}=0$ are highlighted in red. b: Dependence of the $TTS$ together with $L$ on the annealing time $T_{A}$ for the optimal embedding and chain strength combination found in a.
The quality of the sampled solution on the QPU depends strongly on the annealing hyperparameters. We evaluate the quality of a sample set by considering the time to the best solution $TTS$(5) for $p_{S}=0.99$ together with the lowest relative error $L_{min}$ (4) of the sample set. Here, the energy of the best solution found by Gurobi is used as the reference solution. Hence, if $L_{min}=0$, the best solution found by QA is equivalent to the optimal solution found by Gurobi. No optimal solution is obtained if $L_{min}>0$. To generate robust statistics, the sample set for each hyperparameter configuration consists of $N_{s}=10^{5}$ samples.
Similar to the observations made by Gonzalez Calaza et al. [60], we find that the solution quality is sensitive to the used embedding and the chain strength $CS$, that is, the coupling strength of the physical qubits in a chain, see Fig. 4a. For this analysis, we set the annealing time to $T_{A}=50\mu s$ and use the same 10 minor embeddings as in section VI-B. Note that if two embeddings have the same $TTS$ for the same chain strength, only the lower $L_{min}$ is shown Fig. 4a. The full data, showing the performance across all embeddings, is displayed in Fig. 7 in the Appendix. Furthermore, for better comparison between the test cases, we consider the relative chain strength [60] given by
@@
For case 14, the fastest $TTS$ over all ten embeddings is given for $RCS=0.27$; see the left panel in Fig. 4a. The performance of the embeddings varies significantly, especially for small $RCS<0.1$. The optimal solution could not be found for some embeddings and small $RCS$, causing the $TSS$ to reflect only the time needed to find the best solution. For all other configurations, the $TTS$ roughly follows a convex function in the $RCS$. The convex shape indicates that if the chain strength is too small, the chains break frequently, preventing them from uniquely representing a logical bit. On the other hand, if the chain strength is too high, the dynamics during annealing are dominated by the internal dynamics of the chains.
Solving $Q_{part}$ for case 30 to optimality with QA is currently at the limit of what can be achieved within a competitive time frame. c.f. the middle panel in Fig. 4a. For the 8 out of 80 hyperparameter configuration instances where the optimal solution was found, the $TSS$ is $23.03s$, that is, the optimal solution was sampled only once. On the other hand, for most hyperparameter configurations, no optimal solution has been found ($L_{min}>0$), although $L_{min}$ remains relatively small. However, regardless of the $RCS$ and across all embeddings, the best (near-optimal) solutions have only been sampled a few times, manifesting in the uniformly slow $TSS$ values. Inspecting the full data, c.f. Fig. 7 in the Appendix, reveals that the optimal solution could only be found for five out of ten embeddings, highlighting the impact of different embeddings on performance. The best performance is achieved for $RCS=0.15$; optimal solutions have been sampled for four different embeddings. However, the lowest average $L_{min}$ excluding the optimal solutions is achieved for $RCS=0.95$.
For case 57, the current hardware limits are exceeded, at least as a global optimal solver. No optimal solutions could be found for any embedding and chain strength. The lowest average $L_{min}$ is achieved for $RCS=0.15$. Additionally (not shown in the figure), we sampled the QPU $10^{6}$ times for the best-performing combination of $RCS$ and embedding for case 57, which did not significantly improve the solution quality.
Another hyperparameter that, by definition, affects the $TTS$ is the annealing time $T_{A}$. For the best embedding and chain strength combinations from the previous analysis, the $TTS$ increases in $T_{A}$ for all three cases; see Fig. 4b. For case 14, the optimal solution is found for all values of $T_{A}$. Hence, the smallest problems, which do not seem to need extensive hyperparameter optimization, could profit from using the fast annealing feature, which just became available, to achieve an even smaller TTS. On the other hand, for case 30 optimal solutions are only sampled for $T_{A}\geq 25\mu s$. The optimal performance as a global solver is achieved for $T_{A}=25\mu s$. Although for case 57 the optimal solution is never found, the lowest $L_{min}$ is achieved for $T_{A}=50\mu s$, already while scanning for chain strengths, see Fig. 4a. The adiabatic theorem supports the observation that for harder problems, that is problems with a smaller minimum energy gap, longer $T_{A}$ leads to better results since the probability of jumping to a higher energy state throughout the annealing process is reduced.
@@
Gurobi QA
(best hyperparamter)
MIP gap $T_{\text{Gur.}}$ [$s$] $TTS$[$s$] $L_{min}$
Case 14 0 0.43 0.05 0
Case 30 0 4.66 23.03 0
Case 57 0 1.04 23.03 0.01
TABLE I: Comparison of results obtained by Gurobi and QA ignoring all overheads for QA.
We conclude that the performance of QA depends strongly on hyperparameters, particularly on the embedding and the chain strength. Furthermore, for quantum annealing, the actual time to solution suffers from overheads, i. e. the time needed to embed the problem and network latency. For a fair comparison between the performances of quantum annealing on current prototypical hardware with classical deterministic or heuristic solvers, the overhead must be included, and the question arises whether to compare optimized solvers. A fair comparison would thus reveal current hardware limitations of the D-Wave machine, in particular, the need for minor embeddings [62], and is thus not meaningful for the future where some problems might have been overcome, e.g., by adapting the driver to expand the minimal gap [63]. Nevertheless, ignoring all aforementioned overheads and comparing the results obtained by QA for the best hyperparameters with Gurobi, we observe that QA “outperforms” Gurobi running on a laptop only for case 14, see Table I. This observation is consistent with the recent literature: For other highly connected Ising problems, quantum annealing on current D-Wave machines does not provide a speed up or better solution quality [62, 64, 65].